# Testing the Equivalence Principle on Cosmological Scales Using Peculiar Acceleration Power Spectrum

### v1

@@ II.2.1 DM mock generation
4) The covariance matrix for the acceleration power spectrum is accurately estimated using the jackknife method (Norberg et al., 2009) on a single mock dataset. Jackknife realizations are constructed by sequentially deleting one sample at a time and calculating the power spectrum from the remaining data. Specifically, using the delete-one jackknife approach, the simulation box is divided into $n_{s}=64$ equal-volume sub-boxes. Each jackknife box contains $16^{3}$ cells on the CIC mesh. The covariance matrix is estimated as
$C_{ij}=\frac{n_{s}-1}{n_{s}}$ $\sum_{l=1}^{n_{s}}$ $\left[E_{\rm ep}^{[l]}\left(k_{i}\right)-\langle E_{\rm ep}\left(k_{i}\right)\rangle\right]$ (7)
$\times$ $\left[E_{\rm ep}^{[l]}\left(k_{j}\right)-\langle E_{\rm ep}\left(k_{j}\right)\rangle\right]\,,$
@@
II.2.2 Halo mock generation
$z$ $n_{h}$ (low/high) $M_{h}\times 10^{12}~{}M_{\odot}/h$ (low/high)
1.5 $360821/1082464$ $[2.73,3.41]/[3.41,847.050]$
1.0 $466526/1399581$ $[2.73,3.49]/[3.49,1387.33]$
0.5 $536364/1609094$ $[2.73,3.58]/[3.58,2421.35]$
0.0 $550123/1650372$ $[2.73,3.66]/[3.66,5085.91]$
Table 1: Number of halos in the “low-mass” and “high-mass” groups along with their associated mass ranges at different redshifts.
In addition to testing EP using DM particle simulation data, we also conduct the second test by comparing the acceleration power spectra of DM halos with different masses. In the simulation, DM halos are identified using the group finder ROCKSTAR (Behroozi et al., 2013). We selected all halos containing at least 10 DM particles, corresponding to $M_{h}\geq 2.73\times 10^{12}~{}M_{\odot}/h$. These halos are subsequently divided into two mass bins, referred to as “low-mass” and “high-mass”, with the former containing halos of lower mass and the latter containing halos of higher mass. For each fixed redshift, the ratio of the number of low-mass halos to high-mass halos is maintained at approximately $1/3$. This ratio ensures that the shot noise contributions from the small and large mass bins are nearly equal, as confirmed by our measurements. The detailed halo numbers and masses are summarized in Tab. 1. The covariance matrix for $E_{\rm ep}(k)$ is estimated using the same approach described in the previously detailed jackknife method.
@@ III Results
Figure 1: Measurement of $E_{\rm ep}$ for DM particles for testing the Equivalence Principle. The ratio of acceleration power spectra for “group1” and “group2”, each containing $10^{8}$ particles, and the Gaussian redshift error were additionally added in each particle, with zero mean and standard deviation $\sigma_{z_{\parallel}}$, where $f=0.001$ (left) and $f=0.001\times\sqrt{2}$ (right). Four redshifts are considered, and the $k$ values are in the range of $k\in[0.007,0.32]$, with $10$ bins divided on the logarithmic scale. The black solid line indicates the expectation $E_{\rm ep}=1$ if the Equivalence Principle exactly holds. The error bars correspond to the $2\sigma$ level, estimated from the jackknife scheme applied to the simulation box. As seen, in most cases, the $E_{\rm ep}$ values are consistent with unity at the $2\sigma$ level. Note that the bin centers have been slightly adjusted for clarity in the illustration.
$z$ (particle) $\delta_{\rm ep}$ (“A”/“B”) $\sigma_{\delta}$ (“A”/“B”) $\chi^{2}$ (“A”/“B”)
0.0 $0.0004$/$-0.0087$ 0.0098/0.0162 2.71/6.03
0.5 $-0.0097$/$-0.0069$ 0.0081/0.0127 9.17/8.69
1.0 $-0.0040/0.0008$ 0.0064/0.0096 4.05/9.17
1.5 $-0.0018/0.0170$ 0.0057/0.0083 4.50/13.54
Table 2: Best-fit values of $\delta_{\rm ep}$ and the corresponding $1\sigma$ statistical uncertainties, along with the resulting $\chi^{2}$ values for four $z$-bins, are presented for the DM particle simulation data. The results compare two cases (“A” and “B”) with varying SKA observation accuracies.
Figure 2: Same as in Fig. 1, but for the halo mock data, calculated by taking the ratio of acceleration power spectra for the “high-mass” and “low-mass” halo catalogs. In most cases, the derived $E_{\rm ep}$ is consistent with unity within the $2\sigma$ uncertainty level.
$z$ (halo) $\delta_{\rm ep}$ $\sigma_{\delta}$ $\chi^{2}$
0.0 0.0510.051~{}~{}0.0510.051 0.054 4.31
0.5 $-0.021$ 0.038 4.91
1.0 $-0.042$ 0.040 4.09
1.5 $-0.068$ 0.044 12.09
Table 3: Best-fit values of $\delta_{\rm ep}$ and the corresponding $1\sigma$ statistical uncertainties, along with the resulting $\chi^{2}$ values for four $z$-bins, are presented for the DM halo datasets.
Here, we present the measured $E_{\rm ep}$ values for DM particle and halo mock data across four redshift bins ($z=0.0,0.5,1.0,1.5$), with 10 $k$-bins ranging from $0.007$ to $0.32~{}h/\rm Mpc$. The $k$-bins are evenly spaced on a logarithmic scale. For $k\gtrsim 0.3~{}h/\rm Mpc$, the measured power spectra begin to deviate significantly from the linear prediction due to the non-linearity of structure growth (see Appendix A for details). As a result, we restrict our analysis to scales below this threshold.
@@
where correlations between different bins are neglected, as they are negligibly small. The variance $\sigma^{2}_{k}$ in the $k$-th bin is estimated using the jackknife method based on the simulations. Considering that one parameter $\delta$ and 10 $k$-bins are used in the fitting, the degrees of freedom thus are $N_{\rm dof}=10-1=9$.
@@
Fig. 1 presents the measured $E_{\rm ep}$ for DM particle mock data, which serves as a test of the Equivalence Principle. Specifically, we compute the ratio of acceleration power spectra for two groups (as described in Sect. II.2.1). Two different levels of redshift measurement errors are introduced in the simulation particles: the left and right panels correspond to $f=0.001$ and $f=0.001\times\sqrt{2}$, respectively. As shown, the error bars in the right panel are visibly larger than those in the left panel, which is due to the relatively larger $f$ value.
@@
Based on the measured $E_{\rm ep}$, the model fitting results are summarized in Tab. 2, presenting the best-fit values of $\delta_{\rm ep}$, their associated $1\sigma$ statistical uncertainties, and the resulting $\chi^{2}$ values for the four $z$-bins. As seen, the average error in the larger $f$ case is greater than that in the smaller $f$ case by a factor of $1.56$, which is almost consistent with the expected scaling of $\sqrt{2}$. This result demonstrates that the errors are significantly influenced by the redshift measurement error. More importantly, the $E_{\rm ep}$ values are generally consistent with unity within the $2\sigma$ level across most cases. As $N_{\rm dof}=9$, all $\chi^{2}$ values fall within the range of $\chi^{2}\in[0.51,12.48]$, consistent with the $\chi^{2}$ distribution at the $2\sigma$ confidence level. This result validates the $E_{\rm ep}$ estimator as a reliable tool for testing EP at the linear cosmological scales of $k\in[0.007,0.32]~{}h/\rm Mpc$.
@@
In Fig. 2, we present the measured $E_{\rm ep}$ values for the DM halo mock data across four redshift bins. As shown, the $E_{\rm ep}$ values generally agree with the expected values within the $2\sigma$ level. The error bars increase gradually with larger $k$ values, primarily due to the growing influence of shotnoise at smaller scales. After fitting the model, we find that the best-fit $\delta_{\rm ep}$ values are consistent with zero within the $2\sigma$ level, indicating that deviations from unity are not statistically significant. Furthermore, the $\chi^{2}$ values fall within the expected range, supporting the validity of the proposed $E_{\rm ep}$ estimator for the EP test based on the halo simulation data.
@@ IV Conclusion
To achieve this, we develop an EP estimator, $E_{\rm ep}$, designed to evaluate the consistency of peculiar acceleration power spectra across different tracers. If EP holds on cosmological scales for these tracers, the ensemble average of $E_{\rm ep}$ is expected to equal one. As a first step toward testing EP on cosmological scales, we validate this estimator using N-body simulations. Applying this approach, we measure $E_{\rm ep}$ at four redshift bins ($z=[0,0.5,1,1.5]$) on the linear scales of $k\in[0.007,0.32]~{}h/\rm Mpc$. This analysis is conducted using two distinct tests: i) DM particle mock data, incorporating redshift measurement errors, and ii) mock data for low-mass and high-mass DM halos.
To quantitatively assess deviations of the measured $E_{\rm ep}$ from the expected value of unity, we introduce a simple one-parameter model to fit the observed $E_{\rm ep}$. The results indicate that the deviations in different cases approach zero within the $2\sigma$ level, confirming the validity of $E_{\rm ep}$ on the linear cosmological scales.
This study serves as a proof of concept for testing EP using the N-body simulation. Since the simulation do not include any sources that could introduce EP violations, our estimates are expected to confirm the validity of $E_{\rm ep}$ on linear scales, which is indeed verified by our results. In future study, we will apply the $E_{\rm ep}$ estimator to real, high-precision observational data. Additionally, we will incorporate potential EP violations into the simulations to evaluate how effectively the estimator detects such deviations and assess its robustness in analyzing such simulated datasets.
@@
Acknowledgement: This work is supported by National SKA Program of China (2020SKA0110401, 2020SKA0110402, 2020SKA0110100), the National Key R&D Program of China (2020YFC2201600, 2018YFA0404504, 2018YFA0404601), the National Science Foundation of China (12203107, 12473097, 12073088, 12373005, 12273121), the China Manned Space Project with No. CMS-CSST-2021 (A02, A03, B01), the Guangdong Basic and Applied Basic Research Foundation (2024A1515012309). We also wish to acknowledge the Beijing Super Cloud Center (BSCC) and Beijing Beilong Super Cloud Computing Co., Ltd (http://www.blsc.cn/) for providing HPC resources that have significantly contributed to the research results presented in this paper.

### v2

@@ II.3 Measurements of $P_{a_{\parallel}a_{\parallel}}$ and its covariance
The covariance matrix for the acceleration power spectrum or $E_{\rm ep}$ is accurately estimated using the jackknife method (Norberg et al., 2009) on a single mock dataset. Jackknife realizations are constructed by sequentially deleting one sample at a time and calculating the power spectrum from the remaining data. Specifically, using the delete-one jackknife approach, the simulation box is divided into $n_{s}=64$ equal-volume sub-boxes. Each jackknife box contains $16^{3}$ cells on the CIC mesh. The covariance matrix for $E_{\rm ep}$ is estimated as
$C_{ij}=\frac{n_{s}-1}{n_{s}}\sum_{l=1}^{n_{s}}$ $\left[E_{\rm ep}^{[l]}(k_{i})-\langle E_{\rm ep}(k_{i})\rangle\right]$ (9)
$\times$ $\left[E_{\rm ep}^{[l]}(k_{j})-\langle E_{\rm ep}(k_{j})\rangle\right]\,$,
@@ III Results
Here, we present the measured $E_{\rm ep}$ values for DM particle and halo mock data across four redshift bins of $z=[0.0,0.5,1.0,1.5]$, with 10 $k$-bins ranging from $0.007$ to $0.2~h/\rm Mpc$. The $k$-bins are evenly spaced on a logarithmic scale. For $k\gtrsim 0.2~h/\rm Mpc$, the measured power spectrum is roughly an order of magnitude lower than the shot noise level. As a result, we restrict our analysis to scales below this threshold.
@@
where correlations between different bins are neglected, as they are negligibly small. The variance $\sigma^{2}_{k}$ in the $k$-th bin is estimated using the jackknife method based on the simulations. Since the fit involves one free parameter, $\delta_{\rm ep}$, over 10 $k$-bins, the degrees of freedom are $N_{\rm dof}=10-1=9$. The corresponding $2\sigma$ confidence interval for $N_{\rm dof}=9$ corresponds to $\chi^{2}$ values between approximately 2.7 and 19.02.
III.1 Results from DM Particle Mock
Refer to captionRefer to caption
Figure 1: Measurement of $E_{\rm ep}$ for DM particles for testing EP. The ratio of acceleration power spectra for the two datasets–each containing $10^{8}$ particles–was computed after adding Gaussian redshift errors to each particle. These errors have zero mean and a standard deviation of $\sigma_{z_{\parallel}}$, corresponding to $f=0.001$ (left) and $f=0.002$ (right) for the low- and high-noise cases, respectively. Four redshift bins are considered, with $k$ values spanning the range $[0.007,0.21]$ divided into 10 logarithmically spaced bins. The black solid line represents the expected value $E_{\rm ep}=1$ under the exact validity of EP. The error bars correspond to the $1\sigma$ level, estimated from the jackknife scheme applied to the simulation box. The fitted values of $E_{\rm ep}$ are also shown for comparison. As shown, when $f=0.001$, the $E_{\rm ep}$ values remain consistent with unity within the $2\sigma$ confidence level across all $z$-bins. When $f=0.002$, the $E_{\rm ep}$ values exhibit slightly larger deviations from unity and increased uncertainties, particularly at low redshifts. Note that the bin centers have been slightly adjusted for clarity in the illustration.
$z$ (particle) $\delta_{\rm ep}$ ( low/high) $\sigma_{\delta}$ (low/high) $\chi^{2}$ (low/high)
0.0 0.0186/0.1563-0.1563 0.0225/0.0492 9.42/24.09
0.5 0.0084-0.0084/0.0364-0.0364 0.0136/0.0272 9.17/29.59
1.0 0.0108/0.0083 0.0096/0.0184 6.45/8.01
1.5 0.0013/0.0235-0.0235 0.0072/0.0135 5.39/11.08
Table 2: Best-fit values of $\delta_{\rm ep}$ and the corresponding $1\sigma$ statistical uncertainties, along with the resulting $\chi^{2}$ values for four $z$-bins, are presented for the DM particle simulation data. To evaluate the impact of observational noise, the results compare the low- and high-noise cases across different SKA observational accuracies.
@@ III.1 Results from DM Particle Mock
Fig. 1 presents the measured $E_{\rm ep}$ for DM particle mock data, under the fact that EP holds in the underlying physical model. The primary aim here is to assess how redshift measurement uncertainties impact the signal-to-noise of the EP test. Specifically, we compute the ratio of acceleration power spectra for two datasets (as described in Sect. II.2.1). Two different levels of redshift measurement errors are introduced in the simulation particles: the left and right panels correspond to the low-noise case of $f=0.001$ and the high-noise case of $f=0.002$, respectively. As shown, the error bars in the right panel are visibly larger than those in the left panel, which is due to the relatively larger $f$ value.
@@
Based on the measured $E_{\rm ep}$, the model fitting results are summarized in Tab. 2, presenting the best-fit values of $\delta_{\rm ep}$, their associated $1\sigma$ statistical uncertainties, and the resulting $\chi^{2}$ values for the four $z$-bins. As seen, the average error in the high-noise case is greater than that in the low-noise case by a factor of $1.56$, which is almost consistent with the expected scaling of $2$. This result demonstrates that the errors are significantly influenced by the redshift measurement error.
Specifically, in the low-noise case, the fitted values of $\delta_{\rm ep}$ remain close to zero across all redshift bins, with relatively small statistical uncertainties and moderate $\chi^{2}$ values. At $z=0.0$, $\delta_{\rm ep}=0.0186\pm 0.0225$, and the corresponding $\chi^{2}$ is $9.42$, indicating a statistically consistent result with no significant deviation from the null hypothesis. Similar behavior is observed at higher redshifts: at $z=0.5$, the best-fit value is $0.0084\pm 0.0136$ with $\chi^{2}=9.17$; at $z=1.0$, $\delta_{\rm ep}=0.0108\pm 0.0096$ with $\chi^{2}=6.45$; and at $z=1.5$, the value further reduces to $0.0013\pm 0.0072$ with $\chi^{2}=5.39$. These results demonstrate the robustness of the $\delta_{\rm ep}$ estimator in the low-noise case, where statistical errors are well controlled and no significant bias is introduced. In summary, the $E_{\rm ep}$ values are generally consistent with unity within the $2\sigma$ level in most cases. Given the number of degrees of freedom $N_{\rm dof}=9$, all $\chi^{2}$ values lie within the range $\chi^{2}\in[0.51,12.48]$, which is consistent with the expected $\chi^{2}$ distribution at the $2\sigma$ confidence level. These results validate the $E_{\rm ep}$ estimator as a robust and unbiased tool for testing the equivalence principle on linear cosmological scales, specifically over the range $k\in[0.007,0.2]~h/\mathrm{Mpc}$.
In contrast, the high-noise case introduces larger deviations and uncertainties in the fitted $\delta_{\rm ep}$ values, as well as increased $\chi^{2}$ values indicating poorer fit quality. At $z=0.0$, $\delta_{\rm ep}$ shifts to 0.1563-0.1563 with a larger uncertainty of $0.0492$ and a much higher $\chi^{2}=24.09$. At $z=0.5$, the estimate becomes more negative (0.0364-0.0364), and the $\chi^{2}$ rises significantly to $29.59$. These deviations suggest a noise-induced bias that compromises the reliability of the estimator. While the impact of noise is still visible at $z=1.0$ and $z=1.5$, its magnitude decreases: for instance, at $z=1.0$, the best-fit value remains close to zero ($0.0083$), and $\chi^{2}$ increases modestly to $8.01$. At $z=1.5$, $\delta_{\rm ep}=-0.0235$ with $\chi^{2}=11.08$.
Overall, these trends indicate that acceleration noise has a stronger impact at low redshift, leading to biased estimates and reduced statistical significance. In contrast, the low-noise case provides stable and unbiased estimates with tighter constraints. This highlights the critical importance of minimizing observational noise in future SKA surveys, especially for low-redshift measurements aiming to test EP with high precision.
III.2 Results from DM Halo Mock
To evaluate the sensitivity of our estimator to violations of EP, we modified the LOS acceleration by a mass-dependent factor according to Eq. 6, thereby introducing artificial EP-violating effects into the halo catalogs.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 2: Same as Fig. 1, but for the DM halo mock including cases with EP violation. The estimator $E_{\rm ep}$ is computed from the ratio of acceleration power spectra between the “large-mass” and “small-mass” halo catalogs. Shown are the estimated values of the EP violation parameter $\delta_{\rm ep}$ across different redshifts, for various levels of EP violation characterized by the parameters $(\alpha,\beta)$. The case with $(\alpha,\beta)=(0,0)$ shows $\delta_{\rm ep}\approx 0$ across all redshifts, consistent with no violation. Mild violation scenarios such as $(0.1,0.1)$ yield similar results, with $\delta_{\rm ep}$ remaining within $1\sigma$ uncertainties. In contrast, moderate $(0.1,0.3)$ and strong $(0.1,0.5)$ violations produce increasingly significant positive shifts in $\delta_{\rm ep}$, especially at lower redshifts, accompanied by rising $\chi^{2}$ values–indicating robust sensitivity of the estimator to EP-breaking signals.
@@ III.2 Results from DM Halo Mock
In Fig. 2, we present the measured values of the estimator $E_{\rm ep}$ using the DM halo mock data across four redshift bins, considering both EP-preserving and EP-violating scenarios. In the top-left panel, corresponding to the EP-preserving case with $\alpha=0$ and $\beta=0$, the measured $E_{\rm ep}$ values are generally consistent with the expected value of unity within the $2\sigma$ uncertainty range. As the wavenumber $k$ increases, the error bars become larger, primarily due to the increasing contribution of shot noise at smaller scales. After fitting the model, we find that the best-fit $\delta_{\rm ep}$ values are consistent with zero within the $2\sigma$ level, indicating that deviations from unity are not statistically significant. Furthermore, the $\chi^{2}$ values fall within the expected range, supporting the validity of the proposed $E_{\rm ep}$ estimator for the EP test based on the halo simulation data. In particular, as shown in Appendix A, although the halo acceleration power spectrum deviates from the linear theory prediction at $k>0.1~h/{\rm Mpc}$ due to nonlinear structure formation, the close agreement between $P_{a_{\parallel}a_{\parallel}}$ of low- and high-mass halos ensures that the estimator $E_{\rm ep}$ remains consistent with unity up to $k\sim 0.2~h/{\rm Mpc}$. Therefore, the nonlinearity of structure formation does not constitute a significant source of systematic error for our estimator in the relevant scale range.
For the rest of the panels in Fig. 2, we show the results for samples with artificial EP violation, implemented by varying $\beta=0.1$, $0.3$, and $0.5$, as described in Eq. 6. As $\beta$ increases, the deviation of $E_{\rm ep}$ from unity becomes more pronounced. For instance, at redshift $z=0$, the measured $E_{\rm ep}$ values at $k=0.046~h/{\rm Mpc}$ are $1.23$, $1.31$, $1.51$, and $2.15$ for $\beta=0$, $0.1$, $0.3$, and $0.5$, respectively, all with nearly identical uncertainties of $\pm 0.24$. Similar trends are observed at other redshifts, with the magnitude of the deviation increasing approximately proportionally with $\beta$ and becoming more detectable at larger scales (smaller $k$), where shot noise is comparatively lower. These results quantitatively confirm that our estimator $E_{\rm ep}$ is capable of capturing EP-violating signals with high sensitivity over a broad range of redshifts.
Table 3: Summary of best-fit $\delta_{\rm ep}$ values with $1\sigma$ uncertainties $\sigma_{\delta}$ and corresponding $\chi^{2}$ statistics across four redshift bins, for the DM mock catalogs generated under different EP-violation parameters $(\alpha,\beta)$ indicated in the top row.
$\alpha=0$, $\beta=0$
$z$ $\delta_{\rm ep}$ $\sigma_{\delta}$ $\chi^{2}$
0.0 0.0542 0.0587 5.59
0.5 0.0708-0.0708 0.0442 7.72
1.0 0.0485-0.0485 0.0400 4.38
1.5 0.0823-0.0823 0.0476 11.63
$\alpha=0.1$, $\beta=0.1$
$z$ $\delta_{\rm ep}$ $\sigma_{\delta}$ $\chi^{2}$
0.0 0.0853 0.0589 7.20
0.5 0.0503-0.0503 0.0439 6.73
1.0 0.0290-0.0290 0.0402 3.69
1.5 0.0651-0.0651 0.0478 10.97
$\alpha=0.1$, $\beta=0.3$
$z$ $\delta_{\rm ep}$ $\sigma_{\delta}$ $\chi^{2}$
0.0 0.1996 0.0586 18.80
0.5 0.0209 0.0439 6.92
1.0 0.0352 0.0401 5.00
1.5 0.0095-0.0095 0.0475 10.74
$\alpha=0.1$, $\beta=0.5$
$z$ $\delta_{\rm ep}$ $\sigma_{\delta}$ $\chi^{2}$
0.0 0.4366 0.0426 83.87
0.5 0.1562 0.0437 25.37
1.0 0.1597 0.0402 23.80
1.5 0.0968 0.0477 18.88
The results presented in Tab. 3 quantitatively illustrate the sensitivity of the $\delta_{\rm ep}$ estimator to different levels of EP violation, parameterized by $(\alpha,\beta)$. For the no-violation case with $\alpha=0$ and $\beta=0$, the best-fit $\delta_{\rm ep}$ values remain close to zero across all redshifts, ranging from approximately 0.08-0.08 to $0.05$, and are consistent with the $1\sigma$ uncertainties (about 0.040.060.04-0.06). Corresponding $\chi^{2}$ values range from 4.38 to 11.63, indicating statistically acceptable fits.
When a mild violation is introduced ($\alpha=0.1$, $\beta=0.1$), the results remain similar to the no-violation case ($\alpha=0$, $\beta=0$), with $\delta_{\rm ep}$ values exhibiting only slight shifts between 0.065-0.065 and $0.085$ and comparable uncertainties. The $\chi^{2}$ values show no significant increase, indicating that this small violation produces negligible deviation from the null hypothesis. For moderate violation $(\alpha=0.1,\beta=0.3)$, the low-redshift bin ($z=0$) shows a substantial $\delta_{\rm ep}=0.20\pm 0.06$ with $\chi^{2}=18.80$, exceeding the $2\sigma$ threshold and indicating a potential EP violation detection. Higher redshifts exhibit smaller shifts and acceptable $\chi^{2}$ values.
In the strongest violation scenario $(\alpha=0.1,\beta=0.5)$, $\delta_{\rm ep}$ increases markedly, reaching $0.44\pm 0.04$ at $z=0$ with an extreme $\chi^{2}=83.87$, far above the $2\sigma$ limit, strongly rejecting the null hypothesis. The $\delta_{\rm ep}$ values remain significantly above zero at higher redshifts, measuring $0.156\pm 0.043$ at $z=0.5$, $0.159\pm 0.040$ at $z=1.0$, and $0.096\pm 0.047$ at $z=1.5$, with corresponding high $\chi^{2}$ values that indicate notable deviations across these redshifts. Correspondingly, the $\chi^{2}$ values rise sharply up to 83.87 at $z=0$ and remain high at other redshifts, strongly rejecting the no-violation hypothesis. These quantitative trends demonstrate that the estimator robustly detects EP violations, with stronger violations yielding larger $\delta_{\rm ep}$ offsets.
Overall, the observed $\chi^{2}$ values demonstrate that our estimator reliably identifies EP violations, with stronger violations producing larger deviations in both $\delta_{\rm ep}$ and $\chi^{2}$, thus quantitatively confirming the estimator’s effectiveness.
@@ IV Conclusion
To achieve this, we develop an EP estimator, $E_{\rm ep}$, designed to evaluate the consistency of peculiar acceleration power spectra across different tracers. If EP holds on cosmological scales for these tracers, the ensemble average of $E_{\rm ep}$ is expected to equal one. As a first step toward testing EP on cosmological scales, we validate this estimator using N-body simulations. Applying this approach, we measure $E_{\rm ep}$ at four redshift bins ($z=[0,0.5,1,1.5]$) on the linear scales of $k\in[0.007,0.2]~h/\rm Mpc$. This analysis is conducted using two distinct tests: i) DM particle mock data, incorporating redshift measurement errors, and ii) mock data for small-mass and large-mass DM halos.
To quantitatively evaluate deviations of $E_{\rm ep}$ from unity, we fit a one-parameter model to obtain the best-fit $\delta_{\rm ep}$ values across different redshifts. For the DM particle mock assuming no EP violation, we computed the ratio of acceleration power spectra for two DM particle datasets with different redshift measurement errors, characterized by noise parameters $f=0.001$ (low-noise) and $f=0.002$ (high-noise). As expected, the high-noise case exhibits error bars larger by a factor of $\sim 1.56$, close to the theoretical scaling of 2, demonstrating the strong influence of redshift noise on uncertainties. In the low-noise case, the fitted $\delta_{\rm ep}$ values remain close to zero across all redshifts, consistent within the $2\sigma$ confidence interval. Similar results at higher $z$ confirm the estimator’s robustness and absence of bias in low-noise conditions. Conversely, in the high-noise case, $\delta_{\rm ep}$ deviates significantly from zero at low redshifts, indicating noise-induced bias and poorer fit quality. These deviations decrease at higher redshifts as the effect of noise becomes less significant.
Furthermore, for the DM Halo Mock, we considered various EP violation scenarios based on a phenomenological model where the violation strength is controlled by parameters $(\alpha,\beta)$. To quantitatively assess deviations of $E_{\rm ep}$ from unity, we fit the parameter $\delta_{\rm ep}$. For the no-violation case $(\alpha=0,\beta=0)$, $\delta_{\rm ep}$ values are consistent with zero within uncertainties of $\sim 0.04$$0.06$, and $\chi^{2}$ values fall within the $2\sigma$ confidence interval for $N_{\rm dof}=9$. Mild violation $(\alpha=0.1,\beta=0.1)$ produces small shifts in $\delta_{\rm ep}$ and similar $\chi^{2}$, indicating no significant detection. Moderate violation $(\alpha=0.1,\beta=0.3)$ shows a notable deviation at $z=0$, exceeding the $2\sigma$ limit. In the strongest case $(\alpha=0.1,\beta=0.5)$, $\delta_{\rm ep}$ reaches $0.44\pm 0.04$ at $z=0$ with $\chi^{2}=83.87$, and remains significantly positive at higher redshifts with correspondingly high $\chi^{2}$, strongly rejecting the null hypothesis. These results demonstrate the estimator’s robust sensitivity to EP violations, with detection strength scaling with $\beta$.
This study serves as a proof of concept for testing EP using N-body simulations. Since the simulations do not include any physical mechanisms that could induce EP violations, our measurements are expected to confirm the validity of $E_{\rm ep}\approx 1$ on large (linear) scales. This expectation is supported by our results, which show no significant deviation from unity when no artificial EP violation is introduced. Furthermore, we have constructed mock catalogs with artificially introduced EP violations using DM halo samples, and demonstrated that our estimator $E_{\rm ep}$ is capable of detecting such deviations in a statistically significant manner. This confirms that $E_{\rm ep}$ is sensitive to physical violations of EP while remaining robust against shot noise and redshift-space distortions. In future studies, we will apply this estimator to real, high-precision observational data and further refine its application under more complex, observationally realistic conditions, as well as marginalizing over systematic effects to ensure reliable inference. Such efforts will open a new avenue for testing fundamental physics with large-scale structure surveys.
@@
Acknowledgement: We thank Volker Springel for useful discussions. This work is supported by National SKA Program of China (2020SKA0110401, 2020SKA0110402, 2020SKA0110100), the National Key R&D Program of China (2020YFC2201600, 2018YFA0404504, 2018YFA0404601), the National Science Foundation of China (12203107, 12473097, 12073088, 12373005, 12273121), the China Manned Space Project with No. CMS-CSST-2021 (A02, A03, B01), the Guangdong Basic and Applied Basic Research Foundation (2024A1515012309), and the Fundamental Research Funds for the Central Universities, Sun Yat-sen University (No. 24qnpy122). We also wish to acknowledge the Beijing Super Cloud Center (BSCC) and Beijing Beilong Super Cloud Computing Co., Ltd (http://www.blsc.cn/) for providing HPC resources that have significantly contributed to the research results presented in this paper.